# Biologically-Informed Hybrid Membership Inference Attacks on Generative Genomic Models

### v1

@@ 6 Conclusion
Future Directions: Training or finetuning the models on larger more diverse datasets, like using mutations from all chromosomes or leveraging data from large biobanks noauthor_uk_2006 , could lead to better utility. We also believe that evaluating the utility and privacy guarantees of increasingly larger version of MinGPT could unveil the relationship between model utility/privacy and parameter size, leading to some “middle ground model” that offers a natural balance between utility and privacy. Implementing contextual information within the training samples, like population or sex, could further enhance the models’ utility by adding contextual dependencies that would allow the models to better understand the data properties.

### v2

@@ Conclusion
Future Directions
Training or finetuning the models on larger more diverse datasets, like using mutations from all chromosomes or leveraging data from large biobanks (10.1371/journal.pmed.1001779), could lead to better utility. We also believe that evaluating the utility and privacy guarantees of increasingly larger version of MinGPT could unveil the relationship between model utility/privacy and parameter size, leading to some “middle ground model” that offers a natural balance between utility and privacy. Implementing contextual information within the training samples, like population or sex, could further enhance the models’ utility by adding contextual dependencies that would allow the models to better understand the data properties.